OpenStudy (sbentleyaz):

If y = 3x^3 – 2x and dx/dt=3, find dy/dt when x = –2. Give only the numerical answer. For example, if dy/dt = 4, type only 4.

OpenStudy (dumbcow):

\[\frac{dy}{dt} = \frac{dy}{dx}*\frac{dx}{dt}\] \[= (9x^2 -2)(3)\] Plug in x = -2 ...